If (1/x)=(3/4), then what is the value of x?
Answer by algebrahouse.com(1659)

"If (1/x)=(3/4), then what is the value of x?"

1/x = 3/4
3(x) = 4(1) {cross-multiplied}
3x = 4 {multiplied through}
x = 4/3 {divided both sides by 4}
www.algebrahouse.com